zarasht:

At what pressure does ethane (C₂H₆) have a density of 37.2 g/L at 40.0 °C?

justjm:

Use a variation of the ideal gas law \(D = \frac{P*MM}{RT}\) P= you must find R = this is a special constant, it depends on the units of your pressure. If your pressure is asked in atm, then it's 0.082057 L atm 1/mol 1/K T = 40°C, you must convert to K! To do that, °C + 273 = K D = 37.2 g/L MM = go to the periodic table and find the MM of C2H6. Now plug in and solve for P